Which of the following is the solution for x in the simultaneous equations x + y = 24 and 2x – y = -6
1
Expert's answer
2020-08-16T19:32:49-0400

"\begin{alignedat}{2}
 x+ y = 24 \\
 2x-y =-6
\end{alignedat}"

"\begin{alignedat}{2}
 y = 24-x \\
 x+y+2x-y =24-6
\end{alignedat}"




"\begin{alignedat}{2}
 y = 24-x \\
 3x =18
\end{alignedat}"

"\begin{alignedat}{2}
 y = 24-x \\
 x =6
\end{alignedat}"

"\begin{alignedat}{2}
 x=6 \\
 y =18
\end{alignedat}"

"x=6"